### v1

# 3D Diffusion Policy

Abstract
Imitation learning provides an efficient way to teach robots dexterous skills; however, learning complex skills robustly and generalizablely usually consumes large amounts of human demonstrations. To tackle this challenging problem, we present 3D Diffusion Policy (DP3), a novel visual imitation learning approach that incorporates the power of 3D visual representations into diffusion policies, a class of conditional action generative models. The core design of DP3 is the utilization of a compact 3D visual representation, extracted from sparse point clouds with an efficient point encoder. In our experiments involving 72 simulation tasks, DP3 successfully handles most tasks with just 10 demonstrations and surpasses baselines with a 55.3% relative improvement. In 4 real robot tasks, DP3 demonstrates precise control with a high success rate of 85%, given only 40 demonstrations of each task, and shows excellent generalization abilities in diverse aspects, including space, viewpoint, appearance, and instance. Interestingly, in real robot experiments, DP3 rarely violates safety requirements, in contrast to baseline methods which frequently do, necessitating human intervention. Our extensive evaluation highlights the critical importance of 3D representations in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
I Introduction
Imitation learning provides an efficient way to teach robots a wide range of motor skills, such as grasping [68, 61, 81], legged locomotion [42], dexterous manipulation [1, 17, 11], humanoid loco-manipulation [55], and mobile manipulation [58, 13]. Visual imitation learning, which takes high-dimensional visual observations such as images or depth maps, eases the need for task-specific state estimation and thus gains the popularity [10, 61, 81, 12, 21].
However, the generality of visual imitation learning comes at a cost of vast demonstrations [17, 10, 12]. For example, the state-of-the-art method Diffusion Policy [10] necessitates 100 to 200 human-collected demonstrations for each real-world task. To collect the required extensive number of demonstrations, the entire data-gathering process can span several days due to its long-horizon nature and failure-prone process. One solution is online learning [17], where the policy continues to evolve through interaction with environments and a learned reward function from expert demonstrations. Nevertheless, online learning in real-world scenarios introduces its own challenges, such as safety considerations, the necessity for automatic resetting, human intervention, and additional robot hardware costs. Therefore, how to enable (offline) imitation learning algorithms to learn robust and generalizable skills with as few demonstrations as possible is a fundamental problem, especially for practical real-world robot learning.
To tackle this challenging problem, we introduce 3D Diffusion Policy (DP3), a simple yet effective visual imitation learning algorithm that integrates the strengths of 3D visual representations with diffusion policies. DP3 encodes sparsely sampled point clouds into a compact 3D representation using a straightforward and efficient MLP encoder. Subsequently, DP3 denoises random noise into a coherent action sequence, conditioned on this compact 3D representation and the robot poses. This integration leverages not only the spatial understanding capabilities inherent in 3D modalities but also the expressiveness of diffusion models.
To comprehensively evaluate DP3, we have developed a simulation benchmark comprising 72 diverse robotic tasks from 7 domains, alongside a real-world benchmark containing 4 challenging manipulation tasks. Our extensive experiments demonstrate that although DP3 is conceptually straightforward, it exhibits several notable advantages over 2D-based diffusion policies and other baselines:
1.
Efficiency & Effectiveness. DP3 not only achieves superior accuracy but also does so with significantly fewer demonstrations and fewer training steps.
2.
Generalizability. The 3D nature of DP3 facilitates generalization capabilities across multiple aspects: space, viewpoint, instance, and appearance.
3.
Safety. DP3 consistently adheres to all safety requirements in real-world tasks, unlike baseline methods which often give erratic commands and exhibit unexpected behaviors, posing potential damage to the robot hardware.
We conduct several analyses of our 3D visual representations. Intriguingly, we observed that while other baseline methods, such as BCRNN [37] and IBC [12], benefit from the incorporation of 3D representations, they do not achieve enhancements comparable to DP3. Additionally, DP3 consistently outperforms other 3D modalities, including depth and voxel representations, and surpasses other point encoders like PointNeXt [47] and Point Transformer [83]. These ablation studies highlight that the success of DP3 is not just due to the usage of 3D visual representations, but also because of its careful design.
DP3 emphasizes the power of marrying 3D representations with diffusion policies in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
II Related Work
II-A Diffusion Models in Robotics
Diffusion models, a category of generative models that progressively transform random noise into a data sample, have achieved great success in high-fidelity image generation [24, 64, 52, 63]. Owing to their impressive expressiveness, diffusion models have recently been applied in robotics, including in fields such as reinforcement learning [69, 2], imitation learning [10, 41, 51, 71], reward learning [26, 39], grasping [70, 66, 62], and motion planning [53, 28]. In this work, we focus on representing visuomotor policies as conditional diffusion models, referred to as diffusion policies, following the framework established in [10, 41]. Unlike prior methods that primarily focus on images and states as conditions, we pioneer in incorporating 3D conditioning into diffusion policies.
II-B Visual Imitation Learning
Imitation learning offers an efficient way for robots to acquire human-like skills, typically relying on extensive observation-action pairs from expert demonstrations. Given the challenges in accurately estimating object states in the real world, visual observations such as images have emerged as a practical alternative. While 2D image-based policies [40, 12, 10, 37, 17, 57, 68, 16] have predominated the field, the significance of 3D is increasingly recognized [61, 81, 79, 15, 14, 30].
Recent 3D-based policies, including PerAct [61], GNFactor [81], RVT [15], ACT3D [14], NeRFuser [73], and 3D Diffuser Actor [30], have demonstrated notable advancements in low-dimensional control tasks. However, these works face two primary challenges: (1) Impractical setting. These methods convert the imitation learning problem into a prediction-and-planning paradigm using keyframe pose extraction. While effective, this formulation is less suitable for high-dimensional control tasks. (2) Slow inference. The intricate architectures of these methods result in slow inference speeds. For instance, PerAct [61] runs at an inference speed of 2.23 FPS, and 3D Diffuser Actor [30] at 1.67 FPS, making them hard to address tasks that require dense commands, such as highly dynamic environments. Compared to this line of works, we endeavor to develop a universal and fast 3D policy capable of tackling a broader spectrum of robotic tasks, encompassing both high-dimensional and low-dimensional control tasks.
II-C Learning Dexterous Skills
Achieving human-like manipulation skills in robots has been a longstanding objective pursued by robotics researchers. Reinforcement learning has been a key tool in this endeavor, enabling robots with dexterous hands to master a variety of tasks, such as pouring water [48, 80], opening doors [50, 22, 8], rotating objects [45, 75, 77, 46], reorienting objects [19, 7, 6], spinning pens [35], grasping tools [1], executing handovers [82, 25], and building Legos [9]. Imitation learning offers another pathway, with approaches like DIME [3] and DexMV [48] translating human hand movements into robotic actions through retargeting and enabling learning from human videos. Our work, however, diverges from these specific design-centric methods. We demonstrate that enabling the acquisition of these complex skills with minimal demonstrations could be achieved by improving the imitation learning algorithm itself.
III Method
Given a small set of expert demonstrations that contain complex robot skill trajectories, we want to learn a visuomotor policy that maps the visual observations to actions , such that our robots not only reproduce the skill but also generalize beyond the training data. To this end, we introduce 3D Diffusion Policy (DP3), which mainly consists of two critical parts: (a) Perception. DP3 perceives the environments with point cloud data and processes these visual observations with an efficient point encoder into visual features; (b) Decision. DP3 utilizes the expressive Diffusion Policy [10] as the action-making backbone, which generates action sequences conditioning on our 3D visual features. An overview of DP3 is in Figure 2. We will detail each part in the following sections.
III-A A Motivating Example
To better illustrate the generalization ability of DP3, we first give a straightforward example. We use the MetaWorld Reach task [76] as our testbed. In this task, the goal is for the gripper to accurately reach a designated target point. To evaluate the effectiveness of imitation learning algorithms in not only fitting training data but also generalizing to new scenarios, we visualize the training points and the successful evaluation points in 3D space, as shown in Figure 3. We observe that with merely five training points, DP3 reaches points distributed over the 3D space, while for 2D-based methods, Diffusion Policy [10] and IBC [12] learn to reach within a plane-like area, and BCRNN [37] fails to cover the space. This example demonstrates the superior generalization and efficiency of DP3, particularly in scenarios where available data is limited.
III-B Perception
We now detail the perception module in DP3. DP3 focuses on only utilizing a single-view camera for policy learning for all the tasks, which is different from previous works [10, 19] that set up multiple cameras around robots. This is primarily chosen for its practical applicability in real-world tasks.
Representing 3D scenes with point clouds. The 3D scene could be represented in different ways, such as RGB-D images, point clouds, voxels [7], implicit functions [38], and 3D gaussians [31]. Among them, DP3 uses sparse point clouds as the 3D representation. As evidenced in our ablations (see Table IV), point clouds are found to be more efficient compared to other explicit representations, such as RGB-D, depth, and voxels.
For both simulation and the real world, we obtain depth images with size from a single camera. We then convert depth into point clouds with camera extrinsics and intrinsics. We do not use color channels for better appearance generalization.
| Adroit | Bi-DexHands | DexArt | DexDeform | DexMV | HORA | MetaWorld | MetaWorld | MetaWorld | MetaWorld | Average | |
| Algorithm Task | (3) | (6) | (4) | (6) | (2) | (1) | Easy (28) | Medium (11) | Hard (6) | Very Hard (5) | (72) |
| Diffusion Policy | |||||||||||
| DP3 | 87.8 | () |
| Adroit | MetaWorld | DexArt | |||||||||
| Algorithm Task | Hammer | Door | Pen | Assembly | Basketball | Shelf Place | Laptop | Faucet | Toilet | Bucket | Average |
| DP3 | |||||||||||
| Diffusion Policy | |||||||||||
| BCRNN | |||||||||||
| BCRNN+3D | |||||||||||
| IBC | |||||||||||
| IBC+3D | |||||||||||
Point cloud processing. Since the point clouds converted from depth may contain redundant points, such as points from the table and the ground, we crop out these points and only leave points within a bounding box.
We further downsample points by farthest point sampling (FPS, [43]), which helps cover the 3D space sufficiently and reduces the randomness of point cloud sampling, compared to uniform sampling. In practice, we find downsampling or points is sufficient for all the tasks in both simulation and the real world.
Encoding point clouds into compact representations. We then encode point clouds into compact 3D representations with a lightweight MLP network, as shown in Figure 2. The network, termed as DP3 Encoder, is conceptually simple: it consists of a three-layer MLP, a max-pooling function as an order-equivariant operation to pool point cloud features, and a projection head to project the features into a compact vector. LayerNorm layers are interleaved to stabilize training [23]. The final 3D feature, denoted as , is only dimension. As shown in our ablation studies (see Table V), this simple encoder could even outperform pre-trained point encoders such as PointNeXt [47], aligning with observations from [21], where a properly designed small encoder is better than pre-trained large encoders in visuomotor control tasks.
III-C Decision
Conditional action generation. The decision module in DP3 is formulated as a conditional denoising diffusion model [24, 10, 41] that conditions on 3D visual features and robot poses , then denoises a random Gaussian noise into actions . Specifically, starting from a Gaussian noise , the denoising network performs iterations to gradually denoise a random noise into the noise-free action ,
where is Gaussian noise, , , and are functions of and depend on the noise scheduler. This process is also called the reverse process [24].
Training objective. To train the denoising network , we randomly sample a data point from the dataset and do a diffusion process [24] on the data point to get the noise at iteration . The training objective is to predict the noise added to the original data,
where and are noise schedule that performs one step noise adding [24].
Implementation details. We use the convolutional network-based diffusion policy [10]. We use DDIM [63] as the noise scheduler and use sample prediction instead of epsilon prediction for better high-dimensional action generation, with 100 timesteps at training and 10 timesteps at inference. We train 3000 epochs with batch size 128 for DP3 and all the baselines to ensure convergence for all the tasks.
IV Simulation Experiments
IV-A Experiment Setup
Simulation benchmark. Though the simulation environments are increasingly realistic nowadays [36, 72, 65, 84], a notable gap between simulation and real-world scenarios persists [79, 32, 7]. This discrepancy underscores two key aspects: (a) the importance of real robot experiments and (b) the necessity of large-scale diverse simulation tasks for more scientific benchmarking. Therefore, for simulation experiments, we collect in total 72 tasks from 7 domains, covering diverse robotic skills. These tasks range from challenging scenarios like bi-manual manipulation [8], deformable object manipulation [33], and articulated object manipulation [5], to simpler tasks like parallel gripper manipulation [76]. These tasks are built with different simulators including MuJoCo [65], Sapien [72], IsaacGym [36], and PlasticineLab [27], ensuring our benchmarking is not limited by the choice of simulator. Tasks in MetaWorld [76] are categorized into various difficulty levels based on [56]. A brief overview is shown in Table III. The 3D observations are visualized in Figure 4.
Expert demonstrations are collected with agents trained by reinforcement learning (RL) algorithms for all the domains except DexDeform, where we use human-teleoperated data. VRL3 [67] is used for Adroit; BAC [29] is used for MetaWorld; PPO [54] is used in all other domains. We generate successful trajectories with RL agents and ensure all imitation learning algorithms are using the same demonstrations. The success rates for experts are given in Appendix -C.
Baselines. The primary focus of this work is to underscore the significance of the 3D modality in diffusion policies. To this end, our main baseline is the image-based diffusion policy [10], simply referred to as Diffusion Policy. Additionally, we incorporate comparisons with IBC [12], BCRNN [37], and their 3D variations. However, given that these algorithms showed limited effectiveness in our challenging tasks, we evaluate them on only 10 tasks (see Table II). We emphasize that the image and depth resolution for all 2D and 3D methods are the same across all experiments, ensuring a fair comparison.
Evaluation metric. We run 3 seeds for each experiment with seed number . For each seed, we evaluate 20 episodes every 200 training epochs and then compute the average of the highest 5 success rates. We report the mean and std of success rates across 3 seeds.
| Simulation Benchmark (72 Tasks) | ||||||
| Domain | Robo | Object | Simulator | ActD | #Task | #Demo |
| Adroit | Shadow | Rigid/Art | MuJoCo | 28 | 3 | 10 |
| Bi-DexHands | Shadow | Rigid/Art | IsaacGym | 52 | 6 | 10 |
| DexArt | Allegro | Art | Sapien | 22 | 4 | 100 |
| DexDeform | Shadow | Deform | PlasticineLab | 52 | 6 | 10 |
| DexMV | Shadow | Rigid/Fluid | Sapien | 30 | 2 | 10 |
| HORA | Allegro | Rigid | IsaacGym | 16 | 1 | 100 |
| MetaWorld | Gripper | Rigid/Art | MuJoCo | 4 | 50 | 10 |
| Real Robot Benchmark (4 Tasks) | |||||
| Task | Robo | Object | ActD | #Demo | Description |
| Roll-Up | Allegro | Deform | 22 | 40 | Wrap plasticine to make a roll-up |
| Dumpling | Allegro | Deform | 22 | 40 | Wrap plasticine and pinch with fingers |
| Drill | Allegro | Rigid | 22 | 40 | Grasp the drill and touch the cube |
| Pour | Gripper | Rigid | 7 | 40 | Pick the bowl, pour, and place |
IV-B Efficiency and Effectiveness
DP3 shows surprising efficiency across diverse tasks, mainly reflected in the following three perspectives:
1.
High accuracy. Summarized results are in Figure 1(a) and results for each domain are in Table I. We observe that DP3 achieves a success rate exceeding 90% in nearly 30 tasks, whereas Diffusion Policy does in fewer than 15 tasks. Additionally, DP3 did not record any task with a success rate below 10%, in contrast to Diffusion Policy, which had more than 10 tasks below 10%. Note that most of the tasks are only trained with 10 demonstrations.
2.
Learning efficiency. While we train all the algorithms for 3000 epochs to guarantee convergence, we observe that DP3 typically reaches convergence within approximately 500 epochs across all tasks, as illustrated in Figure 5. In contrast, Diffusion Policy tends to converge at a much slower pace or converge into sub-optimal results.
3.
Efficient scaling with demonstrations. As shown in Figure 6, we find that in Adroit tasks, both DP3 and Diffusion Policy perform reasonably while DP3 achieves a comparable accuracy with fewer demonstrations. For MetaWorld tasks above the easy level, Diffusion Policy fails to learn even with an increased number of demonstrations, lagging significantly behind DP3. This underscores that the 3D modality is not just beneficial but essential for certain manipulation tasks.
4.
Competitive inference speed. As depicted in Figure 1, DP3 achieves an inference speed marginally surpassing Diffusion Policy. Contrary to the prevailing assumption that 3D-based policies are slower [61, 81, 71], DP3 manages to achieve efficient inference speeds, primarily attributed to the utilization of sparse point clouds and compact 3D representations.
IV-C Ablations
We select 6 tasks to conduct more ablation studies: Hammer (H), Door (D), Pen (P) from Adroit and Assembly (A), Basketball (B), Shelf Place (S) from MetaWorld. These tasks include both high-dimensional and low-dimensional control tasks, and each task only uses 10 demonstrations. We use the abbreviations of these tasks in the tables for simplicity.
Choice of 3D representations. In DP3, we deliberately select point clouds to represent the 3D scene. To compare different choices of 3D representations, we implement other 3D representations, including RGB-D, depth, and voxel. The RGB-D and depth images are processed using the same image encoder as Diffusion Policy, while voxel representations employ the VoxelCNN, as implemented in [7]. As demonstrated in Table IV, these alternative 3D representations fall short of DP3. Notably, both RGB-D and 2D images are outperformed by depth images, indicating that our tasks are more focused on low-level control rather than scene understanding.
| Repr. | H | D | P | A | B | S | Average |
| Point cloud | |||||||
| Image | |||||||
| Depth | |||||||
| RGB-D | |||||||
| Voxel |
Choice of point cloud encoders. We compare DP3 Encoder with other widely used point encoders, including PointNet [43], PointNet++ [44], PointNeXt [47], and Point Transformer [83]. We also include the pre-trained models of PointNet++ and PointNeXt. Surprisingly, we find that none of these complex models and the pre-trained ones are competitive to DP3 Encoder, as shown in Table V.
| Encoders | H | D | P | A | B | S | Average |
| DP3 Encoder | |||||||
| PointNet | |||||||
| PointNet++ | |||||||
| PointNeXt | |||||||
| Point Transformer | |||||||
| PointNet++ (pre-trained) | |||||||
| PointNeXt (pre-trained) |
Gradually modifying a PointNet. To elucidate the performance disparity between DP3 Encoder and a commonly used point cloud encoder, e.g., PointNet, we gradually modify a PointNet to make it aligned with a DP3 Encoder. Through extensive experiments shown in Table VI, we identify that the T-Net and BatchNorm layers in PointNet are primary inhibitors to its efficiency. By omitting these two elements, PointNet attains an average success rate of , competitive to achieved by our DP3 Encoder. One plausible explanation for the T-Net is that our control tasks use the fixed camera and do not require feature transformations from the T-Net. We would explore the reason for failures of other encoders in the future.
| Encoders | Conv | w/ T-Net | w/ BN | 1024 Dim | Average |
| PointNet | ✓ | ✓ | ✓ | ✓ | |
| Turnaround! | ✓ | ✗ | ✗ | ✓ | |
Design choices in DP3. Besides the 3D representations, the effectiveness of DP3 is contributed by several small design choices, as shown in Table VII. (a) Cropping point clouds helps largely improve accuracy; (b) Incorporating LayerNorm layers could help stabilize training across different tasks [23, 4]; (c) Sample prediction in the noise sampler brings faster convergence, also shown in Figure 7; (d) The projection head in DP3 Encoder accelerates the inference by projecting features to the lower dimension, without hurting accuracy; (e) Removing color channels ensures robust appearance generalization; (f) In low-dimensional control tasks, DPM-solver++ [34] as the noise sampler is competitive to DDIM, while DPM-solver++ could not handle high-dimensional control tasks well.
| Designs | H | D | P | A | B | S | Average |
| DP3 | |||||||
| w/o cropping | |||||||
| w/o LayerNorm | |||||||
| w/o sample pred | |||||||
| w/o projection | |||||||
| w/ color | |||||||
| DDIMDPM-solver++ |
Training policies on demonstrations from different experts. We wanted to study whether the behavior modality of experts would affect the performance of imitation learning methods. To this end, we generate demonstrations with RL-trained agents and script policies respectively in MetaWorld tasks and train on these demonstrations. Summarized results are given in in Table VIII and detailed results are shown in Appendix -C. We could observe that a better expert could generally help train a better policy with imitation learning algorithms, while there still exists a gap between DP3 and the ground truth expert.
| MetaWorld 50 Tasks | Expert (Script) | Expert (BAC [29]) | |
| Ground Truth | |||
| Diffusion Policy | |||
| DP3 |
Different image resolutions for Diffusion Policy. In all experiments, we use the same resolution for DP3 and Diffusion Policy, while we still wanted to study whether improving the image resolution could help Diffusion Policy achieve better results. As shown in Table IX, we observe that a larger resolution such as and could improve Diffusion Policy in the Basketball task, while the success rates averaged across 6 tasks are still close. This again demonstrates the necessity of the 3D representations.
| Algorithm | H | D | P | A | B | S | Average |
| DP3 | |||||||
| Diffusion Policy (84) | |||||||
| Diffusion Policy (128) | |||||||
| Diffusion Policy (256) |
V Real World Experiments
V-A Experiment Setup
Real robot benchmark. DP3 is evaluated across 4 tasks on 2 different robots, including an Allegro hand and a gripper. We use one RealSense L515 camera to obtain real-world visual observations. All the tasks are visualized in Figure 10 and summarized in Table III. Our real-world setup and everyday objects used in our tasks are shown in Figure 8. We now briefly describe our tasks:
1.
Roll-Up. The Allegro hand wraps the plasticine multiple times to make a roll-up.
2.
Dumpling. The Allegro hand first wraps the plasticine and then pinchs it to make dumpling pleats.
3.
Drill. The Allegro hand grasps the drill up and moves towards the green cube to touch the cube with the drill.
4.
Pour. The gripper grasps the bowl, moves towards the plasticine, pours out the dried meat floss in the bowl, and places the bowl on the table.
The randomization in each task is shown in Figure 9. For Roll-Up and Dumpling, the plasticine’s shape and the appearance of the objects placed upon the plasticine are randomized. For Drill and Pour, the variations come from the random positions of the cube, drill, and bowl.
Notably, our tasks using the multi-finger hand are carefully designed to show its advantage over the parallel gripper: In Roll-Up and Dumpling, robots could wrap plasticine without requiring extra tools, unlike RoboCook [60]; In Drill, the drill in the real world is large and heavy, which is quite difficult for the gripper to grasp.
Expert demonstrations are collected by human teleoperation. The Franka arm and the gripper are teleoperated by the keyboard. The Allegro hand is teleoperated with human hands by vision-based retargeting [49, 18]. Since our tasks contain more than one stage and include complex multi-finger robots and deformable objects, making the process of demonstration collection very time-consuming, we only provide 40 demonstrations for each task.
Baselines. Based on our simulation experiments, image-based and depth-based diffusion policies are still powerful, thus we select them as baselines for real-world experiments. Different vision modalities are visualized in Figure 11.
V-B Effectiveness
Results for our real robot tasks are given in Table X. Consistent with our simulation findings, we observe in real-world experiments that DP3 could handle all tasks with high success rates, given only 40 demonstrations. Interestingly, we also observe that while both image-based and depth-based diffusion policies have comparatively low average accuracies, they exhibit distinct strengths in specific tasks. For instance, the image-based diffusion policy excels in the Drill task but fails entirely in Roll-Up. In contrast, the depth-based policy achieves a notable success rate of in Roll-Up.
| Real Robot | Roll-Up | Dumpling | Drill | Pour | Average |
| Diffusion Policy | |||||
| Diffusion Policy (Depth) | |||||
| DP3 |
V-C Generalization
Besides the effectiveness in handling all tasks, DP3 show strong generalization abilities in the real world. We categorize the generalization abilities of DP3 into 4 aspects and detail each aspect as follows.
Spatial generalization. As illustrated in our motivating example, DP3 could better extrapolate in 3D space. We demonstrate this property in the real world, as shown in Table XI. We find that baselines fail to generalize to all test positions while DP3 succeed in 4 out of 5 trials.
![[Uncaptioned image]](x12.png)
| Spatial Generalization | 1 | 2 | 3 | 4 | 5 |
| Diffusion Policy | ✗ | ✗ | ✗ | ✗ | ✗ |
| Diffusion Policy (Depth) | ✗ | ✗ | ✗ | ✗ | ✗ |
| DP3 | ✗ | ✓ | ✓ | ✓ | ✓ |
Appearance generalization. DP3 is designed to process point clouds without color information, inherently enabling it to generalize across various appearances effectively. As demonstrated in Table XII, DP3 consistently exhibits successful generalization to cubes of differing colors, while baseline methods could not achieve. It is noteworthy that the depth-based diffusion policy also does not incorporate color as input. However, due to its lower accuracy on the training object, the ability to generalize is also limited.
One solution to improve the appearance generalization ability of image-based methods is applying strong data augmentation during training [21, 20], which however could impede the learning process [78, 20]. More importantly, the primary objective of this work is to demonstrate that DP3, even without the aid of any data augmentation, can effectively generalize, thereby underscoring the potential of 3D representations in real robot learning.
| Apperance Generalization () | |||||
| Diffusion Policy | ✗ | ✗ | ✗ | ✗ | ✗ |
| Diffusion Policy (Depth) | ✗ | ✗ | ✗ | ✗ | ✗ |
| DP3 | ✓ | ✓ | ✓ | ✓ | ✓ |
Instance generalization. Achieving generalization across diverse instances, which vary in shape, size, and appearance, presents a significantly greater challenge compared to mere appearance generalization. In Table XIII, we demonstrate that DP3 effectively manages a wide range of everyday objects. This success can be primarily attributed to the inherent characteristics of point clouds. Specifically, the use of point clouds allows for policies that are less prone to confusion, particularly when these point clouds are downsampled. This feature significantly enhances the model’s ability to adapt to varied instances.
![[Uncaptioned image]](x14.png)
| Instance Generalization | Mouse | Espeon | Cup | Mug | Hand |
| Diffusion Policy | ✗ | ✗ | ✗ | ✗ | ✓ |
| Diffusion Policy (Depth) | ✗ | ✗ | ✓ | ✗ | ✗ |
| DP3 | ✓ | ✓ | ✓ | ✓ | ✓ |
View generalization. Generalizing image-based methods across different views is notably challenging [74], and acquiring training data from multiple views can be time-consuming and costly [81, 59]. We demonstrate in Table XIV that DP3 effectively addresses this generalization problem when the camera views are altered slightly. It is important to note that since the camera view is altered, we manually adjust the cropped space of the point clouds.
![[Uncaptioned image]](x15.png)
| View Generalization | View 1 | View 2 | View 3 |
| Diffusion Policy | ✗ | ✗ | ✗ |
| Diffusion Policy (Depth) | ✗ | ✗ | ✗ |
| DP3 | ✓ | ✓ | ✓ |
V-D Safety Violation
In our real-world experiments, we surprisingly observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors, which necessitates human termination to ensure robot safety. This situation is defined as safety violation. We compute the safety violation rate in our main real-world experiments, shown in Table XV. Interestingly, DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning.
![[Uncaptioned image]](x16.png)
| Safety Violation Rate | Roll-Up | Dumpling | Drill | Pour | Average |
| Diffusion Policy | |||||
| Diffusion Policy (Depth) | |||||
| DP3 |
VI Conclusion
In this work, we introduce 3D Diffusion Policy (DP3), an efficient visual imitation learning algorithm, adept at managing a wide range of robotic tasks in both simulated and real-world environments with only a small set of demonstrations. The essence of DP3 lies in its integration of carefully designed 3D representations with the expressiveness of diffusion policies. Across 72 simulated tasks, DP3 outperforms its 2D counterpart by a relative margin of . In real-world scenarios, DP3 shows high accuracy in executing complex manipulations of deformable objects using the Allegro hand. More importantly, we demonstrate that DP3 possesses robust generalization capabilities across various aspects and causes fewer safety violations in real-world scenarios.
Limitations. Though we have developed an efficient architecture, the optimal 3D representation for control is still yet discovered. Besides, this work does not delve into tasks with extremely long horizons, which remains for future exploration.
Acknowledgement
We would like to thank Zhecheng Yuan, Cheng Lu, and Jianfei Chen for their helpful discussions. This work is supported by National Key R&D Program of China (2022ZD0161700).